nikiwe moves into her new flat and buys furniture worth R7500. the furniture company charges finance rates of 20% per year for 3 years, at simple interest rate. calculate the interest over 3 years
Answer by stanbon(75887)

nikiwe moves into her new flat and buys furniture worth R7500. the furniture company charges finance rates of 20% per year for 3 years, at simple interest rate. calculate the interest over 3 years
--------
I = P*r*t
-----
I = 7500*0.20*3 = R4500
----------------
Cheers,
Stan H.
---------------